The cost of direct materials transferred into the Bottling Department of the Mountain Springs Water Company is $650,000. The conversion cost for the period in the Bottling Department is $498,000. The total equivalent units for direct materials and conversion are 26,000 liters and 8,300 liters, respectively. Determine the direct materials and conversion cost per equivalent unit.

Cost per equivalent unit of materials $
Cost per equivalent unit of conversion $
Answer #1

Answer

The following is the formula for finding cost per equivalent unit

Cost per equivalent unit = cost / equivalent units

Direct material cost is $650,000

Conversion cost is $498,000

Equivalent unit of material = 26,000 liters (unit)

Equivalent unit of conversion cost = 8,300 liters

Cost per equivalent unit of material

= 650,000 / 26,000 = $25

Cost per equivalent unit of conversion cost

= 498,000 / 8,300 = $60

cost per equivalent unit of material is $25 and cost per equivalent unit of conversion cost is $60.

If YOU are like the answer please upvote.